# Mitigating Suboptimality of Deterministic Policy Gradients in Complex Q-functions

### v1

@@ Abstract
In reinforcement learning, off-policy actor-critic approaches like DDPG and TD3 are based on the deterministic policy gradient. Herein, the Q-function is trained from off-policy environment data and the actor (policy) is trained to maximize the Q-function via gradient ascent. We observe that in complex tasks like dexterous manipulation and restricted locomotion, the Q-value is a complex function of action, having several local optima or discontinuities. This poses a challenge for gradient ascent to traverse and makes the actor prone to get stuck at local optima. To address this, we introduce a new actor architecture that combines two simple insights: (i) use multiple actors and evaluate the Q-value maximizing action, and (ii) learn surrogates to the Q-function that are simpler to optimize with gradient-based methods. We evaluate tasks such as restricted locomotion, dexterous manipulation, and large discrete-action space recommender systems and show that our actor finds optimal actions more frequently and outperforms alternate actor architectures.
@@ 1 Introduction
In sequential decision-making, the goal is to build an optimal agent that maximizes the expected cumulative returns (Sondik, 1971; Littman, 1996). Value-based reinforcement learning (RL) approaches learn each action’s expected returns with a Q-function and maximize it (Sutton & Barto, 1998). However, in continuous action spaces, evaluating the Q-value of every possible action is impractical. This necessitates an actor to globally maximize the Q-function and efficiently navigate the vast action space (Grondman et al., 2012). Yet, this is particularly challenging in tasks such as restricted locomotion, with a non-convex Q-function landscape with many local optima (Figure 2).
Figure 2: Complex Q-landscapes. We plot Q-value versus action $a$ for some state. In control of Inverted-Double-Pendulum-Restricted (left) and Hopper-Restricted (middle), certain action ranges are unsafe, resulting in various locally optimal action peaks. In a large discrete-action recommendation system (right), there are local peaks at actions representing real items (black dots).
Can we build an actor architecture to find closer to optimal actions in such complex Q-landscapes? Prior methods perform a search over the action space with evolutionary algorithms like CEM (De Boer et al., 2005; Kalashnikov et al., 2018; Shao et al., 2022), but this requires numerous costly re-evaluations of the Q-function. To avoid this, deterministic policy gradient (DPG) algorithms (Silver et al., 2014), such as DDPG (Lillicrap et al., 2015) and TD3 (Fujimoto et al., 2018) train a parameterized actor to output actions with the objective of maximizing the Q-function locally.
@@
To improve actors’ ability to identify optimal actions in complex, non-convex Q-function landscapes, we propose the Successive Actors for Value Optimization (SAVO) algorithm. SAVO leverages two key insights: (1) combining multiple policies using an $\arg\max$ on their Q-values to construct a superior policy (§4.1), and (2) simplifying the Q-landscape by excluding lower Q-value regions based on high-performing actions, inspired by tabu search (Glover, 1990), thereby reducing local optima and facilitating gradient-ascent (§4.2). By iteratively applying these strategies through a sequence of simplified Q-landscapes and corresponding actors, SAVO progressively finds more optimal actions.
We evaluate SAVO in complex Q-landscapes such as (i) continuous control in dexterous manipulation (Rajeswaran et al., 2017) and restricted locomotion (Todorov et al., 2012), and (ii) discrete decision-making in the large action spaces of simulated (Ie et al., 2019) and real-data recommender systems (Harper & Konstan, 2015), and gridworld mining expedition (Chevalier-Boisvert et al., 2018). We use the reframing of large discrete action RL to continuous action RL following Van Hasselt & Wiering (2009) and Dulac-Arnold et al. (2015), where a policy acts in continuous actions, such as the feature space of recommender items (Figure 2), and the nearest discrete action is executed.
Our key contribution is SAVO, an actor architecture to find better optimal actions in complex non-convex Q-landscapes (§4). In experiments, we visualize how SAVO’s successively learned Q-landscapes have fewer local optima (§6.2), making it more likely to find better action optima with gradient ascent. This enables SAVO to outperform alternative actor architectures, such as sampling more action candidates (Dulac-Arnold et al., 2015) and learning an ensemble of actors (Osband et al., 2016)6.1) across continuous and discrete action RL.
@@ 3 Problem Formulation
Our work tackles the effective optimization of the Q-value landscape in off-policy actor-critic methods for continuous and large-discrete action RL. We model a task as a Markov Decision Process (MDP), defined by a tuple $\{\mathcal{S},\mathcal{A},\mathcal{T},R,\gamma\}$ of states, actions, transition probabilities, reward function, and a discount factor. The action space $\mathcal{A}\!\subseteq\!\mathbb{R}^{D}$ is a $D$-dimensional continuous vector space. At every step $t$ in the episode, the agent receives a state observation $s_{t}\in\mathcal{S}$ from the environment and acts with $a_{t}\in\mathcal{A}$. Then, it receives the new state after transition $s_{t+1}$ and a reward $r_{t}$. The objective of the agent is to learn a policy $\pi(a\mid s)$ that maximizes the expected discounted reward, $\max_{\pi}\mathbb{E}_{\pi}\left[\sum_{t}\gamma^{t}r_{t}\right].$
@@ 3.1 Deterministic Policy Gradients (DPG)
DPG (Silver et al., 2014) is an off-policy actor-critic algorithm that trains a deterministic actor $\mu_{\phi}$ to maximize the Q-function. This happens via two steps of generalized policy iteration, GPI (Sutton & Barto, 1998): policy evaluation estimates the Q-function (Bellman, 1966) and policy improvement greedily maximizes the Q-function. To approximate the $\arg\max$ over continuous actions in Eq. (2), DPG proposes the policy gradient to update the actor locally in the direction of increasing Q-value,
@@
$\nabla_{\phi}J(\phi)$ $=\mathbb{E}_{s\sim\rho^{\mu}}\left[\nabla_{a}Q^{\mu}(s,a)\big{|}_{a=\mu(s)}\nabla_{\phi}\mu_{\phi}(s)\right].\vspace{-15pt}$ (3)
@@ 3.2 The Challenge of an Actor Maximizing a Complex Q-landscape
DPG-based algorithms train the actor following the chain rule in Eq. (3). Specifically, its first term, $\nabla_{a}Q^{\mu}(s,a)$ involves gradient ascent in Q-versus-$a$ landscape. This Q-landscape is often highly non-convex (Figures 23) and non-stationary because of its own training. This makes the actor’s output $\mu(s)$ get stuck at suboptimal Q-values, thus leading to insufficient policy improvement in Eq. (2). We can define the suboptimality of the $\mu$ w.r.t. $Q^{\mu}$ at state $s$ as
@@
Suboptimality in actors is a crucial problem because it leads to (i) poor sample efficiency by slowing down GPI, and (ii) poor inference performance even with an optimal Q-function, $Q^{*}$ as seen in Figure 3 where a TD3 actor gets stuck at a locally optimum action $a_{0}$ in the final Q-function.
Figure 3: Non-convex Q-landscape in Inverted-Pendulum-Restricted leads to a suboptimally converged actor.
@@
In contrast, our challenge involves non-convexity in the Q-function w.r.t. the action space. The actor’s task is to find, for every state $s$, the action $a$ that maximizes $Q^{\mu}(s,a)$. Since the Q-function can be highly non-convex and multimodal in $a$, the gradient ascent step $\nabla_{a}Q^{\mu}(s,a)$ used in Eq. (3) may lead the actor to converge to suboptimal local maxima in action space. Unlike parameter space optimization, the actor cannot rely on high dimensionality or overparameterization to smooth out the optimization landscape in action space because the Q-landscape is determined by the task’s reward. Furthermore, the non-stationarity of the Q-function during training compounds this challenge. These properties make our non-convex challenge unique, requiring a specialized actor to navigate the complex Q-landscape.
Tasks with several local optima in the Q-function include restricted inverted pendulum shown in Figure 3, where certain regions of the action space are invalid or unsafe, leading to a rugged Q-landscape (Florence et al., 2022). Dexterous manipulation tasks exhibit discontinuous behaviors like inserting a precise peg in place with a small region of high-valued actions (Rajeswaran et al., 2017) and surgical robotics have a high variance in Q-values of nearby motions (Barnoy et al., 2021).
@@ 3.2.1 Large Discrete Action RL Reframed as Continuous Action RL
We discuss another practical domain where non-convex Q-functions are present. In large discrete action tasks like recommender systems (Zhao et al., 2018; Zou et al., 2019; Wu et al., 2017), a common approach (Van Hasselt & Wiering, 2009; Dulac-Arnold et al., 2015) is to use continuous representations of actions as a medium of decision-making. Given a set of actions, =$\mathcal{I}=\{\mathscr{I}_{1},\dots,\mathscr{I}_{N}\}$, a predefined module $\mathcal{R}:\mathcal{I}\rightarrow\mathcal{A}$ assigns each $\mathscr{I}\in\mathcal{I}$ to its representation $\mathcal{R}(\mathscr{I})$, e.g., text embedding of a given movie (Zhou et al., 2010). A continuous action policy $\pi(a\mid s)$ is learned in the action representation space, with each $a\in\mathcal{A}$ converted to a discrete action $\mathscr{I}\in\mathcal{I}$ via nearest neighbor,
@@
Importantly, the nearest neighbor operation creates a challenging piece-wise continuous Q-function with suboptima at various discrete points as shown in Figure 2 (Jain et al., 2021; 2020).
@@ 4 Approach: Successive Actors for Value Optimization (SAVO)
Our objective is to design an actor architecture that efficiently discovers better actions in complex, non-convex Q-function landscapes. We focus on gradient-based actors and introduce two key ideas:
@@
    Maximizing Over Multiple Policies: By combining policies using an $\arg\max$ over their Q-values, we can construct a policy that performs at least as well as any individual policy (§4.1).
@@
    Simplifying the Q-Landscape: Drawing inspiration from tabu search (Glover, 1990), we propose using actions with good Q-values to eliminate or “tabu” the Q-function regions with lower Q-values, thereby reducing local optima and facilitating gradient-based optimization (§4.2).
4.1 Maximizer Actor over Action Proposals
@@
Here, $\mu_{M}$ is shown to be a better maximizer of $Q(s,a)$ in Eq. (2) than $\mu\;\forall s$ :
@@
Therefore, by policy improvement theorem (Sutton & Barto, 1998), $V^{\mu_{M}}(s)\geq V^{\mu}(s)$, proving that $\mu_{M}$ is better than a single $\mu$ for a given $Q$. Appendix A proves the following theorem by showing that policy evaluation and improvement with $\mu_{M}$ converge.
@@
A modified policy iteration algorithm where $\nu_{0}=\mu$ is the current policy learned with DPG and maximizer actor $\mu_{M}$ defined in Eq. (5), converges in the tabular setting to the optimal policy.
@@
4.2 Successive Surrogates to Reduce Local Optima
To train additional policies $\nu_{i}$ that can improve upon $\mu_{M}$, we introduce surrogate Q-functions with fewer local optima, inspired by the principles of tabu search (Glover, 1990), which is an optimization technique that uses memory structures to avoid revisiting previously explored inferior solutions, thereby enhancing the search for optimal solutions. Similarly, our surrogate functions act as memory mechanisms that “tabu” certain regions of the Q-function landscape deemed suboptimal based on previously identified good actions. Given a known action $a^{\dagger}$, we define a surrogate function that elevates the Q-values of all inferior actions to $Q(s,a^{\dagger})$, which serves as a constant threshold:
@@ 4.2 Successive Surrogates to Reduce Local Optima
Extending this idea, we define a sequence of surrogate functions using the actions from previous policies. Let $a_{<i}=\{a_{0},a_{1},\dots,a_{i-1}\}$. The $i$-th surrogate function is:
@@
Proof Sketch. As $\Psi_{i}\!\to\Psi_{i+1}$, the anchor Q-value in Eq. (7) weakly increases, $\max_{j<i}Q(s,a_{j})\leq\max_{j<(i+1)}Q(s,a_{j})$, thus, eliminating more local minima below it (proof in Appendix B.1). ∎
4.3 Successive Actors for Surrogate Optimization
To effectively reduce local optima using the surrogates $\Psi_{1},\dots,\Psi_{k}$, we design the policies $\nu_{i}$ to optimize their respective surrogates $\Psi_{i}(s,a;a_{<i})$. Each $\nu_{i}$ focuses on regions where $Q(s,a)\geq\max_{j<i}Q(s,a_{j})$, allowing it to find better optima than previous policies. The actor $\nu_{i}$ is conditioned on previous actions $\{a_{0},\dots,a_{i-1}\}$, summarized using deep sets (Zaheer et al., 2017) (Figure 4). The maximizer actor $\mu_{M}$ (Eq. (5)) selects the best action among all proposals.
We train each actor $\nu_{i}$ using gradient ascent on its surrogate $\Psi_{i}$, similar to DPG:
$\nabla_{\phi_{i}}J(\phi_{i})=\mathbb{E}_{s\sim\rho^{\mu_{M}}}\left[\nabla_{a}\Psi_{i}(s,a;a_{<i})\big{|}_{a}\,\nabla_{\phi_{i}}\nu_{i}(s;a_{<i})\right].$ (8)
4.4 Approximate Surrogate Functions
The surrogates $\Psi_{i}$ have zero gradients when $Q(s,a)<\tau$, where $\tau=\max_{j<i}Q(s,a_{j})$,
$\nabla_{a}\Psi_{i}(s,a;a_{<i})=\begin{cases}\nabla_{a}Q^{\mu_{M}}(s,a)&\text{if }Q(s,a)\geq\tau,\\ 0&\text{if }Q(s,a)<\tau.\end{cases}$
This means the policy gradient only updates $\nu_{i}$ when $Q(s,a)\geq\tau$, which may slow down learning. To address this issue, we ease the gradient flow by learning a smooth lossy approximation $\hat{\Psi}_{i}$ of $\Psi_{i}$.
Figure 5: In restricted inverted pendulum, given an anchor $Q(a_{0})$ value, $\Psi$ (left) has some zero-gradient surfaces which $\hat{\Psi}$ (right) approximately follows while allowing non-zero gradients towards high Q-values to flow into its actor $\nu$.
We approximate each surrogate $\Psi_{i}$ with a neural network $\hat{\Psi}_{i}$. This approach leverages the universal approximation theorem (Hornik et al., 1989; Cybenko, 1989) and benefits from empirical evidence that deep networks can effectively learn non-smooth functions (Imaizumi & Fukumizu, 2019). The smooth surrogate $\hat{\Psi}_{i}$ enables continuous gradient propagation, which is essential for optimizing the actors $\nu_{i}$. We train $\hat{\Psi}_{i}$ to approach $\Psi_{i}$ by minimizing the mean squared error at two critical points:

### v2

@@ Abstract
In reinforcement learning, off-policy actor-critic methods like DDPG and TD3 use deterministic policy gradients: the Q-function is learned from environment data, while the actor maximizes it via gradient ascent. We observe that in complex tasks such as dexterous manipulation and restricted locomotion with mobility constraints, the Q-function exhibits many local optima, making gradient ascent prone to getting stuck. To address this, we introduce SAVO, an actor architecture that (i) generates multiple action proposals and selects the one with the highest Q-value, and (ii) approximates the Q-function repeatedly by truncating poor local optima to guide gradient ascent more effectively. We evaluate tasks such as restricted locomotion, dexterous manipulation, and large discrete-action space recommender systems and show that our actor finds optimal actions more frequently and outperforms alternate actor architectures.
@@ 1 Introduction
In sequential decision-making, the goal is to build an optimal agent that maximizes the expected cumulative returns (Sondik, 1971; Littman, 1996). Value-based reinforcement learning (RL) approaches estimate the future returns of an action with a Q value, then select actions that maximize this Q value (Sutton & Barto, 1998). In continuous action spaces, directly enumerating all actions is impractical, so an actor is introduced to learn which actions yield the maximum Q-value (Grondman et al., 2012). We show that common continuous control benchmarks (Lillicrap et al., 2015) exhibit easily optimized Q functions, which obscures a key challenge in current RL algorithms. Specifically, when the Q-function is non-convex, such as locomotion with restricted mobility in Figure 1, a learning actor can produce suboptimal behavior by converging at one of the local optima.
Figure 1: In continuous control tasks, we visualize trained TD3 Q-values at a fixed state $s_{f}$ over the full action space (projected to 2D), $Q(s,a|s=s_{f})$. In Inverted-Double-Pendulum (left) and Hopper (middle) with action ranges restricted due to immobility, Q-landscapes have multiple local optima. In a large discrete-action recommendation task (right), local peaks correspond to real items (black dots). In such non-convex Q-landscapes, gradient-based actors often converge to suboptimal actions.
Can we build an actor architecture to find better optimal actions in such complex Q-landscapes? Prior methods perform a search over the action space with evolutionary algorithms like CEM (De Boer et al., 2005; Kalashnikov et al., 2018; Shao et al., 2022), but this requires numerous costly re-evaluations of the Q-function. To avoid this, deterministic policy gradient (DPG) algorithms (Silver et al., 2014), such as DDPG (Lillicrap et al., 2015), TD3 (Fujimoto et al., 2018), and REDQ (Chen et al., 2020) train a parameterized actor to output actions with the objective of maximizing the Q-function locally.
@@
To improve actors’ ability to identify optimal actions in complex, non-convex Q-function landscapes, we propose the Successive Actors for Value Optimization (SAVO) algorithm. SAVO leverages two key insights: (1) combining multiple policies using an $\arg\max$ on their Q-values to construct a superior policy (§4.1), and (2) simplifying the Q-landscape by excluding lower Q-value regions based on high-performing actions, inspired by Tabu Search (Glover, 1990)—a metaheuristic that avoids cycling back to recently visited suboptimal solutions by maintaining an explicit memory of them. SAVO achieves this via a sequence of surrogate Q-functions that iteratively exclude the Q-value of regions below previously identified inferior actions, thereby reducing local optima and facilitating gradient ascent (§4.2), enabling the corresponding actors to discover higher-quality actions.
We evaluate SAVO in complex Q-landscapes such as (i) continuous control in dexterous manipulation (Rajeswaran et al., 2017) and restricted locomotion (Todorov et al., 2012), and (ii) discrete decision-making in the large action spaces of simulated (Ie et al., 2019) and real-data recommender systems (Harper & Konstan, 2015), and gridworld mining expedition (Chevalier-Boisvert et al., 2018). We use the reframing of large discrete action RL to continuous action RL following (Van Hasselt & Wiering, 2009) and Dulac-Arnold et al. (2015), where a policy acts in continuous actions, such as the feature space of recommender items (Figure 1), and the nearest discrete action is executed.
Our key contribution is SAVO, an actor architecture to find better optimal actions in complex non-convex Q-landscapes (§4). In experiments, we visualize how SAVO’s successively learned Q-landscapes have fewer local optima (§6.2, §18), making it more likely to find better action optima with gradient ascent. This enables SAVO to outperform alternative actor architectures, such as sampling more action candidates (Dulac-Arnold et al., 2015) and learning an ensemble of actors (Osband et al., 2016)6.1) across continuous and discrete action RL.
@@ 3 Problem Formulation
Our work tackles the effective optimization of the Q-value landscape in off-policy actor-critic methods for continuous and large-discrete action RL. We model a task as a Markov Decision Process (MDP), defined by a tuple $\{\mathcal{S},\mathcal{A},\mathcal{T},R,\gamma\}$ of states, actions, transition probabilities, reward function, and a discount factor. The action space $\mathcal{A}$ is a $D$-dimensional continuous vector space, $\mathbb{R}^{D}$. At every step $t$ in the episode, the agent receives a state observation $s_{t}\in\mathcal{S}$ from the environment and acts with $a_{t}\in\mathcal{A}$. Then, it receives the new state after transition $s_{t+1}$ and a reward $r_{t}$. The objective of the agent is to learn a policy $\pi(a\mid s)$ that maximizes the expected discounted reward, $\max_{\pi}\mathbb{E}_{\pi}\left[\sum_{t}\gamma^{t}r_{t}\right].$
@@ 3.1 Deterministic Policy Gradients (DPG)
DPG (Silver et al., 2014) is an off-policy actor-critic algorithm that trains a deterministic actor $\mu_{\phi}$ to maximize the Q-function. This happens via two steps of generalized policy iteration, GPI (Sutton & Barto, 1998): policy evaluation estimates the Q-function (Bellman, 1966) and policy improvement greedily maximizes the Q-function. To approximate the $\arg\max$ over continuous actions in Eq. 2, DPG proposes the policy gradient to update the actor locally in the direction of increasing Q-value,
@@
$\nabla_{\phi}J(\phi)$ $=\mathbb{E}_{s\sim\rho^{\mu}}\left[\nabla_{a}Q^{\mu}(s,a)\big|_{a=\mu(s)}\nabla_{\phi}\mu_{\phi}(s)\right].\vskip-15.0pt$ (3)
@@ 3.2 The Challenge of an Actor Maximizing a Complex Q-landscape
DPG-based algorithms train the actor following the chain rule in Eq. 3. Specifically, its first term, $\nabla_{a}Q^{\mu}(s,a)$ involves gradient ascent in Q-versus-$a$ landscape. This Q-landscape is often highly non-convex (Fig. 13) and changes non-stationarily during training. This makes the actor’s output $\mu(s)$ get stuck at suboptimal Q-values, thus leading to insufficient policy improvement in Eq. 2. We can define the suboptimality of the $\mu$ w.r.t. $Q^{\mu}$ at state $s$ as
@@
Figure 3: Non-convex Q-landscape in Inverted-Pendulum-Restricted leads to the TD3 actor converging at a local optimum $a_{0}$ with large suboptimality.
Suboptimality in actors is a crucial problem because it leads to (i) poor sample efficiency by slowing down GPI, and (ii) poor inference performance even with an optimal Q-function, $Q^{*}$ as seen in Fig. 3 where a TD3 actor gets stuck at a locally optimum action $a_{0}$ in the final Q-function.
@@
In contrast, our challenge involves non-convexity in the Q-function w.r.t. the action space. The actor’s task is to find, for every state $s$, the action $a$ that maximizes $Q^{\mu}(s,a)$. Since the Q-function can be highly non-convex and multimodal in $a$, the gradient ascent step $\nabla_{a}Q^{\mu}(s,a)$ used in Eq. 3 may lead the actor to converge to suboptimal local maxima in action space. Unlike parameter space optimization, the actor cannot rely on high dimensionality or overparameterization to smooth out the optimization landscape in action space because the Q-landscape is determined by the task’s reward. Furthermore, the non-stationarity of the Q-function during training compounds this challenge. These properties make our non-convex challenge unique, requiring a specialized actor to navigate the complex Q-landscape.
Tasks with several local optima in the Q-function include inverted pendulum with restricted action space due to limitations on mobility or terrain, leading to a rugged Q-landscape (Florence et al., 2022) as shown in Fig. 3. Dexterous manipulation tasks exhibit discontinuous behaviors like inserting a precise peg in place with a small region of high-valued actions (Rajeswaran et al., 2017) and surgical robotics have a high variance in Q-values of nearby motions (Barnoy et al., 2021).
@@ 3.2.1 Large Discrete Action RL Reframed as Continuous Action RL
We discuss another practical domain where non-convex Q-functions are present. In large discrete action tasks like recommender systems (Zhao et al., 2018; Zou et al., 2019; Wu et al., 2017), a common approach (Van Hasselt & Wiering, 2009; Dulac-Arnold et al., 2015) is to use continuous representations of actions as a medium of decision-making. Given a set of actions, $\mathcal{I}=\{\mathscr{I}_{1},\dots,\mathscr{I}_{N}\}$, a predefined module $\mathcal{R}:\mathcal{I}\rightarrow\mathcal{A}$ assigns each $\mathscr{I}\in\mathcal{I}$ to its representation $\mathcal{R}(\mathscr{I})$, e.g., text embedding of a given movie (Zhou et al., 2010). A continuous action policy $\pi(a\mid s)$ is learned in the action representation space, with each $a\in\mathcal{A}$ converted to a discrete action $\mathscr{I}\in\mathcal{I}$ via nearest neighbor,
@@
Importantly, the nearest neighbor operation creates a challenging piece-wise continuous Q-function with suboptima at various discrete points as shown in Fig. 1 (Jain et al., 2021; 2020).
@@ 4 Approach: Successive Actors for Value Optimization (SAVO)
We propose an online actor architecture and training method that dynamically guides gradient-based policy improvement toward better actions throughout training. Our method preserves the time-efficiency of gradient-based methods as opposed to maximization using expensive evolutionary methods while mitigating the suboptimality of a single actor. We introduce two key ideas:
@@
    Multiple Actors: We train several gradient-based actors and select among their proposed actions via $\arg\max$ on the Q-function, ensuring the resulting policy outperforms any single actor (§4.1).
@@
    Easier to maximize Q-landscape: We train online surrogates of the Q-function that are biased towards higher-value actions and progressively flatten out shallow local maxima so that gradient-based improvement is likely to find actions in better regions (§4.2).
While surrogates generate diverse candidate actions, the final decision always uses an $\arg\max$ over the true Q-function estimate, ensuring we never do worse than ignoring the surrogates altogether.
4.1 Maximizer Actor over Multiple Action Proposals
@@
$\mu_{M}$ can be simply shown to be a better maximizer of $Q(s,a)$ in Eq. 2 than $\mu\;\forall s$ :
@@
Therefore, by policy improvement theorem (Sutton & Barto, 1998), $V^{\mu_{M}}(s)\geq V^{\mu}(s)$, proving that $\mu_{M}$ is better than a single $\mu$ for a given $Q$. Appendix 9 proves the following theorem by showing that policy evaluation and improvement with $\mu_{M}$ converge.
@@
A modified policy iteration algorithm where $\nu_{0}=\mu$ is the current policy learned with DPG and maximizer actor $\mu_{M}$ defined in Eq. 5, converges in the tabular setting to the locally optimal policy.
@@
4.2 Successive Q-landscape surrogates for Better Action Proposals
To obtain better-than-random action proposals for $\mu_{M}$, we train additional policies $\nu_{i}$ with gradient-ascent on surrogate Q-functions with three properties:
  1. 1.
    Truncate regions below anchor actions: We train online surrogates of the Q-function that are biased towards higher-value actions and progressively flatten out shallow local maxima so that gradient-based improvement is likely to find actions in better regions.
  2. 2.
    Approximately track Q-function with a bias towards high valued actions: We train several gradient-based actors and select among their proposed actions via $\arg\max$ on the Q-function, ensuring the resulting policy outperforms any single actor.
  3. 3.
    Gradient-based actors for each surrogate: Each surrogate likely provides a path to progressively better optima for its actor, which in turn provides a better anchor for the following surrogates.
4.2.1 Truncate regions below anchor actions
Our inspiration is Tabu Search (Glover & Laguna, 1998), which is an optimization technique that avoids revisiting previously explored inferior solutions, thereby enhancing the search for optimal solutions. We adopt a similar idea by identifying relatively high-value actions, referred to as anchor actions, which serve as local reference points during optimization. Specifically, we propose to “tabu” regions of the Q-function landscape that fall below these anchors’ values. Concretely, we define a surrogate function $\Psi$ that truncates the landscape by elevating the Q-values of all inferior actions to $Q(s,a^{\dagger})$, effectively flattening suboptimal basins and guiding the actor away from poor local optima,
@@ 4.2.1 Truncate regions below anchor actions
Extending this idea, we define a sequence of surrogate functions using the actions from all previous policies as anchors. Let $a_{<i}=\{a_{0},a_{1},\dots,a_{i-1}\}$ be the anchors, the $i$-th surrogate function is:
@@
Proof Sketch. As $\Psi_{i}\!\to\Psi_{i+1}$, the anchor Q-value in Eq. 7 weakly increases, $\max_{j<i}Q(s,a_{j})\leq\max_{j<(i+1)}Q(s,a_{j})$, thus, eliminating more local minima below it (proof in Appendix 10.1). ∎
4.2.2 Approximately track Q-function with a bias towards high valued actions
The surrogates $\Psi_{i}$ have zero gradients in the flattened regions when $Q(s,a)<\tau$, where $\tau=\max_{j<i}Q(s,a_{j})$, This means the policy gradient only updates $\nu_{i}$ when $Q(s,a)\geq\tau$, which may slow down learning. To address this issue, we ease the gradient flow by learning a smooth approximation $\hat{\Psi}_{i}$ of $\Psi_{i}$, that is biased towards high-valued actions to provide a path to a better optimum.
Figure 5: While $\Psi$ (left) has flat surfaces, $\hat{\Psi}$ (right) smoothens the function to allow non-zero gradients to flow into the actor towards better optima in Inverted-Pendulum-Restricted.
We approximate each surrogate $\Psi_{i}$ with a neural network $\hat{\Psi}_{i}$, by training it with imitation learning to track the updates to the Q-function (that is being updated by TD error) at two critical actions:
$\mathcal{L}_{\text{approx}}=\mathbb{E}_{s\sim\rho^{\mu_{M}}}\left[\sum_{a\in\{\tilde{\mu}_{M}(s),\nu_{i}(s;a_{<i})\}}\left\|\hat{\Psi}_{i}(s,a;a_{<i})-\Psi_{i}(s,a;a_{<i})\right\|_{2}^{2}\right],\text{where}$ (8)